OpenStudy (anonymous):

Hi. I have a question about a problem. It is: Calculate the derivative of each function directly from the definition. f(x) equals 1/(2x+1). The answers say it is -2/(2x+1)squared. I understand everything except how to get the correct denominator. Any help would be great!

OpenStudy (anonymous):

whew... by the definition...

OpenStudy (anonymous):

When they say directly from the definition, they mean evaluate the following: \[ \lim_{h\to0}\frac{f(x+h)-f(x)}{h}. \]So, we plug in \(f(x)\), and obtain:\[ \lim_{h\to0}\frac{\frac{1}{2(x+h)+1}-\frac{1}{2x+1}}{h}. \]Now we simplify this:\[ \lim_{h\to0}\frac{\frac{1}{2x+2h+1}-\frac{1}{2x+1}}{h}=\lim_{h\to0}\frac{\frac{2x+1-(2x+2h+1)}{(2x+2h+1)(2x+1)}}{h}=\lim_{h\to0}\frac{\frac{-2h}{(2x+2h+1)(2x+1)}}{h}\\ =\lim_{h\to0}\frac{-2}{(2x+2h+1)(2x+1)}=\frac{-2}{(2x+1)(2x+1)}=\frac{-2}{(2x+1)^2}\]

OpenStudy (anonymous):

Sorry, had made a typo. Do you follow?

OpenStudy (anonymous):

Yeah except I do not understand how the denominator simplifies to (2x+1(2x+1).

OpenStudy (anonymous):

just plug in h=0 you end up with (2x+0+1) = (2x+1)

OpenStudy (anonymous):

Oh h=0...I kept trying to multiply it out and it wasn't giving me the right answer (obviously). I guess I just spaced out. Thanks I appreciate it!